I am having trouble with this homework problem:
North Carolina State University posts the grade distributions for its courses online. 5 Students in Statistics 101 in the Fall 2007 semester received 26% A’s, 42% B’s, 20% C’s, 10% D’s, and 2% F’s. Choose a Statistics 101 student at random. To “choose at random” means to give every student the same chance to be chosen. The student’s grade on a four-point scale (with A = 4) is a discrete random variable X with this probability distribution:
x=0 probability=0.02
x=1 probability=0.10
x=2 probability=0.20
x=3 probability=0.42
x=4 probability=0.26
(a) Say in words what the meaning of P(X ≥ 3) is. What is this probability?
(b) Write the event “the student got a grade poorer than C” in terms of values of the random variable X. What is the probability of this event?
For (a) I said: if P(x>_3) then P is either less than or equal to 3 and this means that P is .42+.26 all added together to reach the probability of .68 or 68% chance the students grade is at least a B or better.
For (b) I said that if the student got a grade poorer than C than P(x<2)or P(x Is this correct?



Answer by ewatrrr(24785)


Hi,
x = 4 , 3 , 2 , 1, 0 represents an A,B,C,D or F grade
P(X ≥ 3)
(a) Say in words what the meaning of P(X ≥ 3) is.
What is this probability of a student receiving a grade of B or better in Stat 101
P(X ≥ 3)= P(3)+ P(4) = .42 + .26 = .68 Or 68%
(b) Write the event “the student got a grade poorer than C” in terms of values of the random variable X. What is the probability of this event?
(X < 2) , P(X <2 ) = P(1) + P(0) = .10 + .02 = .12 Or 12 %